# Phys4DGen: Physics-Compliant 4D Generation with Multi-Material Composition Perception

### v4

@@ Abstract
4D content generation aims to create dynamically evolving 3D content that responds to specific input objects such as images or 3D representations. Current approaches typically incorporate physical priors to animate 3D representations, but these methods suffer from significant limitations: they not only require users lacking physics expertise to manually specify material properties but also struggle to effectively handle the generation of multi-material composite objects. To address these challenges, we propose Phys4DGen, a novel 4D generation framework that integrates multi-material composition perception with physical simulation. The framework achieves automated, physically plausible 4D generation through three innovative modules: first, the 3D Material Grouping module partitions heterogeneous material regions on 3D representations surfaces via semantic segmentation; second, the Internal Physical Structure Discovery module constructs the mechanical structure of object interiors; finally, we distill physical prior knowledge from multimodal large language models to enable rapid and automatic material properties identification for both objects’ surfaces and interiors. Experiments on both synthetic and real-world datasets demonstrate that Phys4DGen can generate high-fidelity 4D content with physical realism in open-world scenarios, significantly outperforming state-of-the-art methods.
Figure 1: The red arrows indicate the direction of external forces. We use the space-time slice (right column), where the vertical axis represents time and the horizontal axis show a spatial slice of the object (marked by red lines), to reveal motion intensity and frequency. As shown in top, diffusion models embed unrealistic motion priors that may mislead the estimation process—e.g., Physics3D consistently overestimates the softness of the ficus, deviating from physical plausibility. Additionally, the accuracy of such approaches heavily depends on the setting of initial material properties (e.g., Init Youngs’modulus $10^{4}$ vs. $10^{6}$). In contrast, our method achieves more accurate material properties estimation within 14.88 seconds, enabling reliable simulation.
@@ 1 Introduction
The creation of 4D content has gained substantial importance across multiple domains including computer animation, interactive gaming, and immersive virtual reality applications [4]. In particular, recent advances in generative models have fundamentally transformed 4D generation due to their powerful visual priors [32, 47, 39, 5, 26, 38]. They leverage dynamic priors from video diffusion models [49, 30, 46, 45, 35], enabling the automated production of high-quality 4D content. However, these data-driven approaches fundamentally lack physical modeling constraints, often resulting in generated motions that violate physical laws and exhibit noticeable inconsistencies.
@@ 2 Related Work
4D generation aims to generate dynamic 3D content that aligns with user input conditions such as text, images, and videos. Unlike 3D generation [23, 27, 36, 19, 40, 16], which primarily focuses on producing spatially consistent geometry and appearance, 4D generation must additionally ensure temporal realism and consistency across frames, making the task significantly more challenging. Based on the input conditions, 4D generation can be categorized into three types: text-to-4D [50, 3, 2, 33, 18], video-to-4D [44, 8, 11, 25, 46], and image-to-4D [49, 43, 35, 42]. MAV3D  [33] first employs temporal SDS from the text-to-video diffusion model to optimize HexPlane [7] representation. 4D-fy [3] introduces hybrid score distillation during training for high-quality text-to-4D generation. Instead of text input, Consistent4D [13] uses SDS for geometry optimization and interpolation loss for spatiotemporal consistency in 4D generation from monocular video. 4DGen [44] supervised by pseudo labels generated from multi-view diffusion model. Animate124 [49] pioneered an image-to-4D framework using a coarse-to-fine strategy that combines different diffusion priors. Generating 4D content from an image in DreamGaussian4D  [30] avoids using temporal SDS and instead performs optimization based on reference videos generated by a video diffusion model. Our framework can generates physically plausible and temporally coherent 4D content efficiently, without extra optimization steps.
Figure 2: Framework of Phys4DGen. (a) 3D Gaussians Generation: Given an input image, a static 3D Gaussians is generated under the guidance of the diffusion model. (b) Material Grouping and Internal Discover: 3D Material Grouping is applied to partition the 3D Gaussians into distinct material groups. Concurrently, Internal Physical Structure Discover is used to fill internal particles and determine their corresponding material groups. (c) MLLMs-Guided Material Identification: Surface and internal material properties are visually inferred by MLLMs. These inferred results are then integrated into the 3D representation $\mathbb{P}$ through the CLIP Fusion module, forming a material continuum representation $\tilde{\mathbb{P}}$. (d) 4D Dynamics Generation: Given external forces, MPM simulation is performed to animate the material continuum, thereby generating 4D content.
@@ 2.2 Physics-Based Dynamic Generation
Some recent methods have attempted to leverage physical simulation to create visually dynamics with physical realism. PhysGen [22] introduces rigid-body physics simulation to achieve physically grounded video synthesis from a single image. PAC-NeRF [17] recovers object geometry and physical properties from multi-view videos by combining NeRF with differentiable physics, without requiring known shapes. However, NeRF’s implicit representation is not ideal for physical simulation. The explicit representation of 3DGS [14] through a set of anisotropic Gaussian kernels, which can be interpreted as particles in space, enables the many applications of physical simulations [41, 10, 51]. PhysGaussian [41] is the first to apply MPM to 3D Gaussian representations, enabling the simulation of realistic physical dynamics. Spring-Mass [51] introduces a novel integration of a spring-mass system into the 3DGS framework for elastic material simulation. However, they require manual specification of material types, material properties, and simulation regions. PhysDreamer [48] utilizes reference videos from video diffusion models for supervision to estimate material properties. Likewise, DreamPhysics and Physics3D [12, 20] use score distillation sampling from video diffusion models to optimize physical properties. The stochastic nature of video diffusion models and their non-physical dynamic priors can distort material property estimation. Our method does not require an optimization process and can efficiently infer material properties. Notably, we uniquely consider the possibility that a single object may comprise multiple materials, including differing internal and external compositions.
@@ 3 Methodology
The proposed Phys4DGen is illustrated in Fig 2. Given a single input image, a static 3D Gaussians representation is generated under the guidance of a diffusion model (Sec. 3.1). We then employ 3D Material Grouping (Sec. 3.2) to assign suitable material groups to different regions of the 3D Gaussians. To further explore internal details, Internal Physical Structure Discovery (Sec. 3.3) is employed to model internal geometries and assign the internal material groupings. Subsequently, MLLM-Guided Material Identification (Sec. 3.4) infer both surface and internal material properties for each material region from the input image. These are fused into the 3D representation through CLIP Fusion, producing a material continuum representation. Finally, given external forces and boundary conditions, the 4D dynamics are simulated using MPM (Sec. 3.5) based on the material continuum.
@@ 3.1 Static 3D Gaussians Generation
Given the recent advances in image-to-3D generation [38, 37], which have demonstrated the ability to produce high-quality 3D content, we directly employ these methods to generate static 3D Gaussians, providing a plug-and-play capability within our framework. We choose 3DGS as a representation for its explicit representation nature. 3DGS represents 3D objects using a collection of anisotropic Gaussian kernels [14], which can be interpreted as particles in space. Thus, 3D Gaussians can be viewed as a discretization of the continuum, which is highly beneficial for integrating particle-based physical simulation algorithms. In this phase, we obtain a static 3D Gaussians $\boldsymbol{\mathcal{G}}$ for subsequent simulation. Each Gaussian kernel can be represented as $\boldsymbol{\mathcal{G}_{k}}=(\mathbf{x}_{k},\mathbf{\Sigma}_{k},\alpha_{k},\mathbf{c}_{k})$. Here, $\mathbf{x}_{k}$, $\mathbf{\Sigma}_{k}$, $\alpha_{k}$, and $\mathbf{c}_{k}$ represent the center position, covariance matrix, opacity, and color of the Gaussian kernel $k$, respectively. Additionally, the quality of the 4D content improves with the quality of the static 3D Gaussians generated from the input image.
@@
Given the static 3D Gaussians generated in the previous phase, we render the scene from multiple views, producing an image sequence and its corresponding depth maps. SAM2 [15, 29] is a powerful vision foundation model, supports accurate video segmentation. To ensure consistency of 2D mask maps across views—i.e., that the same material region receives the same mask index from different views— we consider the multi-view image sequence as a video input and apply SAM2 for segmentation to generate the associate mask maps $\mathbf{M}=\{\mathbf{M}_{o}\}_{o=1}^{N}$.
@@
where $\mathbf{K}$ and $[\mathbf{R}_{o}|\mathbf{T}_{o}]$ represent the camera’s intrinsic and extrinsic parameters, respectively. We use the 3DGS-estimated depth to check if the Gaussian kernel is visible in the segmentation map $\mathbf{M}_{o}$. if it is visible, we include the material group from this view in the voting process. After processing the segmentation maps for all views, we perform majority voting, assigning the material group $g_{k}$ that appears most frequently across all views to Gaussian kernel $\boldsymbol{\mathcal{G}_{k}}={(\mathbf{x}_{k},\mathbf{\Sigma}_{k},\alpha_{k},\mathbf{c}_{k},g_{k})}$. Repeating these steps allows us to determine the material groups for all Gaussian kernels $\mathcal{G}$.
@@ 3.3 Physical Internal Structure Discovery
Since the internal structure of an object can only be inferred from surface information available in the input image, we establish a surface-to-interior correspondence by assigning material groups from surface particles to internal ones. To ensure the closure of the boundary set, we additionally designate the outermost layer of grid particles as surface particles. Each surface particle is then assigned to its nearest Gaussian kernel via a nearest-neighbor search and inherits its material group accordingly. For each internal particle, rays are cast along the six principal axes to collect material groups from the intersected surface particles. The most frequently observed material group $g_{i}$ is then assigned to the internal particle. Subsequently, we merge 3D Gaussians $\mathcal{G}$ with the filled internal particles $\mathcal{P}_{I}$ to obtain a unified continuum particles representation with material groups: =$\mathbb{P}=\{\mathcal{G}|\mathcal{P}_{I}\}=\{\{(\mathbf{x}_{k},g_{k})\}|\{(\mathbf{x}_{i},g_{i})\}\}$. Based on this representation, we can easily assign material properties to different material regions of the 3D object.
@@ 3.4.1 Material Information Reasoning
In the real world, objects are typically composed of different materials, In 4D generation, users often lack the necessary physical knowledge to provide reasonable material properties for simulation. This greatly limits the physical realism of the simulation results. Recently, multimodal large language models have advanced rapidly, exhibiting knowledge far beyond that of humans, including rich physical prior knowledge. Inspired by this, we introduce GPT-4o [1], which reasons the material properties (e.g., Density $\rho$, Young’s modulus $E$, Poisson’s ratio $\nu$) of the internal and external parts of objects through vision.
Specifically, the user-input image is treated as the canonical view $\mathbf{I}_{c}$ for reasoning with the MLLMs. Sub-images representing different material components are extracted from this view based on the segmentation mask map. Following this, the canonical view and its segmented sub-images are fed into GPT-4o, prompting it to reason the internal and external material types and properties for the object described in each segmented sub-image. Detailed prompts are provided in Appendix. The output format is as follows: {partial object 1: {surface: {material type, material property}, internal: {material type, material property}}, …}.
3.4.2 CLIP Fusion
@@
Next, we calculate the similarity between the CLIP features of the canonical view and those of the rendered image. By selecting the matching pairs with the highest similarity, we establish a correspondence between the material groupings. Based on this, we can assign the surface and internal material information to the corresponding continuum particles. Finally, we construct a complete material continuum representation ~$\tilde{\mathbb{P}}={(\mathbf{x}_{p},g_{p},\theta_{p})}$, where $\theta_{p}$ denotes the material properties, for use in subsequent physical simulation to generate 4D dynamics.
@@
Phys4DGen can integrate any particle-based physical simulation algorithm. In this paper, we use the Material Point Method (MPM) [41] to simulate the dynamics of 4D content, which enables the modeling of motion and deformation behavior of continuum under external forces. For details on MPM and external forces, please refer to Appendix. For physical simulation, we further assign temporal properties $t$ to the material continuum, along with other physical attributes involved in the simulation process, such as mass $m$, deformation gradient $\mathbf{F}$, and velocity $\mathbf{v}$. Then, we employ MPM to perform physical simulations on the material continuum ~$\tilde{\mathbb{P}}^{t}$. This allows us to track the position and local deformation of each particle at every time step:
@@
where $\mathbf{x}^{t+1}=\{\mathbf{x}_{p}^{t+1}\}_{p=1}^{P}$ denotes the positions of all particles at time step $t+1$. $\mathbf{F}^{t+1}=\{\mathbf{F}_{p}^{t+1}\}_{p=1}^{P}$ represents deformation gradients, which describe the local deformation of each particle at time step $t+1$. To reconstruct the 3D Gaussian Splatting (3DGS) representation at time step $t$ from the simulation results, we isolate the 3DGS-relevant components from the simulated material continuum. To incorporate the local deformation behavior of each GS kernel, we interpret the deformation gradient as a local affine transformation applied to the Gaussian kernel. Consequently, we can derive the covariance matrix of Gaussian kernel $k$ in step $t+1$:
@@
At each step of the MPM simulation, we obtain the deformed 3DGS representation. The sequence of 3DGS representations across all time steps collectively forms the 4D content. This enables the generation of physcally plausible 4D dynamics. Furthermore, since material grouping for 3D objects has already been established in Sec. 3.2, external forces can be precisely applied to any specific material region without manual selection, enabling user-friendly interaction and fine-grained control.

### v5

@@ Abstract
4D content generation aims to create dynamically evolving 3D content that responds to specific input objects such as images or 3D representations. Current approaches typically incorporate physical priors to animate 3D representations, but these methods suffer from significant limitations: they not only require users lacking physics expertise to manually specify material properties but also struggle to effectively handle the generation of multi-material composite objects. To address these challenges, we propose Phys4DGen, a novel 4D generation framework that integrates multi-material composition perception with physical simulation. The framework achieves automated, physically plausible 4D generation through three innovative modules: first, the 3D Material Grouping module partitions heterogeneous material regions on 3D representations’ surfaces via semantic segmentation; second, the Internal Physical Structure Discovery module constructs the mechanical structure of object interiors; finally, we distill physical prior knowledge from multimodal large language models to enable rapid and automatic material properties identification for both objects’ surfaces and interiors. Experiments on both synthetic and real-world datasets demonstrate that Phys4DGen can generate high-fidelity 4D content with physical realism in open-world scenarios, significantly outperforming state-of-the-art methods.
Figure 1: The red arrows indicate the direction of external forces. We use the space-time slice (right column), where the vertical axis represents time and the horizontal axis shows a spatial slice of the object (marked by red lines), to reveal motion intensity and frequency. As shown in the top, diffusion models embed unrealistic motion priors that may mislead the estimation process—e.g., Physics3D consistently overestimates the softness of the ficus, deviating from physical plausibility. Additionally, the accuracy of such approaches heavily depends on the setting of initial material properties (e.g., Init Young’s modulus $10^{4}$ vs. $10^{6}$). In contrast, our method achieves more accurate material properties estimation within 14.88 seconds, enabling reliable simulation.
@@ 1 Introduction
The creation of 4D content has gained substantial importance across multiple domains, including computer animation, interactive gaming, and immersive virtual reality applications [4]. In particular, recent advances in generative models have fundamentally transformed 4D generation due to their powerful visual priors [38, 54, 45, 6, 31, 44]. They leverage dynamic priors from video diffusion models [56, 36, 53, 52, 41], enabling the automated production of high-quality 4D content. However, these data-driven approaches fundamentally lack physical modeling constraints, often resulting in generated motions that violate physical laws and exhibit noticeable inconsistencies.
@@ 2 Related Work
4D generation aims to generate dynamic 3D content that aligns with user input conditions such as text, images, and videos. Unlike 3D generation [27, 32, 42, 21, 46, 17], which primarily focuses on producing spatially consistent geometry and appearance, 4D generation must additionally ensure temporal realism and consistency across frames, making the task significantly more challenging. Based on the input conditions, 4D generation can be categorized into three types: text-to-4D [57, 3, 2, 39, 20], video-to-4D [51, 9, 12, 30, 53], and image-to-4D [56, 50, 41, 49]. MAV3D  [39] first employs temporal SDS from the text-to-video diffusion model to optimize HexPlane [8] representation. 4D-fy [3] introduces hybrid score distillation during training for high-quality text-to-4D generation. Instead of text input, Consistent4D [14] uses SDS for geometry optimization and interpolation loss for spatiotemporal consistency in 4D generation from monocular video. 4DGen [51] supervised by pseudo labels generated from multi-view diffusion model. Animate124 [56] pioneered an image-to-4D framework using a coarse-to-fine strategy that combines different diffusion priors [25, 26]. Generating 4D content from an image in DreamGaussian4D  [36] avoids using temporal SDS and instead performs optimization based on reference videos generated by a video diffusion model. Our framework can generate physically plausible and temporally coherent 4D content efficiently, without extra optimization steps.
Figure 2: Framework of Phys4DGen. (a) 3D Gaussians Generation: Given an input image, a static 3D Gaussians is generated under the guidance of the diffusion model. (b) Material Grouping and Internal Discover: 3D Material Grouping is applied to partition the 3D Gaussians into distinct material groups. Concurrently, Internal Physical Structure Discovery is used to fill internal particles and determine their corresponding material groups. (c) MLLMs-Guided Material Identification: Surface and internal material properties are visually inferred by MLLMs. These inferred results are then integrated into the 3D representation $\mathbb{P}$ through the CLIP Fusion module, forming a material continuum representation $\tilde{\mathbb{P}}$. (d) 4D Dynamics Generation: Given external forces, MPM simulation is performed to animate the material continuum, thereby generating 4D content.
@@ 2.2 Physics-Based Dynamic Generation
Some recent methods have attempted to leverage physical simulation to create visual dynamics with physical realism. PhysGen [24] introduces rigid-body physics simulation to achieve physically grounded video synthesis from a single image. PAC-NeRF [18] recovers object geometry and physical properties from multi-view videos by combining NeRF with differentiable physics, without requiring known shapes. However, NeRF’s implicit representation is not ideal for physical simulation. The explicit representation of 3DGS [15, 47] through a set of anisotropic Gaussian kernels, which can be interpreted as particles in space, enables the many applications of physical simulations [48, 11, 58, 19]. PhysGaussian [48] is the first to apply MPM to 3D Gaussian representations, enabling the simulation of realistic physical dynamics. Spring-Mass [58] introduces a novel integration of a spring-mass system into the 3DGS framework for elastic material simulation. However, they require manual specification of material types, material properties, and simulation regions. PhysDreamer [55] utilizes reference videos from video diffusion models for supervision to estimate material properties. Likewise, DreamPhysics and Physics3D [13, 22] use score distillation sampling from video diffusion models to optimize physical properties. The stochastic nature of video diffusion models and their non-physical dynamic priors can distort material property estimation. Our method does not require an optimization process and can efficiently infer material properties. Notably, we uniquely consider the possibility that a single object may comprise multiple materials, including differing internal and external compositions.
@@ 3 Methodology
The proposed Phys4DGen is illustrated in Fig 2. Given a single input image, a static 3D Gaussians representation is generated under the guidance of a diffusion model (Sec. 3.1). We then employ 3D Material Grouping (Sec. 3.2) to assign suitable material groups to different regions of the 3D Gaussians. To further explore internal details, Internal Physical Structure Discovery (Sec. 3.3) is employed to model internal geometries and assign the internal material groupings. Subsequently, MLLM-Guided Material Identification (Sec. 3.4) infers both surface and internal material properties for each material region from the input image. These are fused into the 3D representation through CLIP Fusion, producing a material continuum representation. Finally, given external forces and boundary conditions, the 4D dynamics are simulated using MPM (Sec. 3.5) based on the material continuum.
@@ 3.1 Static 3D Gaussians Generation
Given the recent advances in image-to-3D generation [44, 43], which have demonstrated the ability to produce high-quality 3D content, we directly employ these methods to generate static 3D Gaussians. We choose 3DGS as a representation for its explicit representation nature. 3DGS represents 3D objects using a collection of anisotropic Gaussian kernels [15], which can be interpreted as particles in space. Thus, 3D Gaussians can be viewed as a discretization of the continuum, which is highly beneficial for integrating particle-based physical simulation algorithms. In this phase, we obtain a static 3D Gaussians $\boldsymbol{\mathcal{G}}$ for subsequent simulation. Each Gaussian kernel can be represented as $\boldsymbol{\mathcal{G}_{k}}=(\mathbf{x}_{k},\mathbf{\Sigma}_{k},\alpha_{k},\mathbf{c}_{k})$. Here, $\mathbf{x}_{k}$, $\mathbf{\Sigma}_{k}$, $\alpha_{k}$, and $\mathbf{c}_{k}$ represent the center position, covariance matrix, opacity, and color of the Gaussian kernel $k$, respectively. Additionally, the quality of the 4D content improves with the quality of the static 3D Gaussians generated from the input image.
@@
Given the static 3D Gaussians generated in the previous phase, we render the scene from multiple views, producing an image sequence and its corresponding depth maps. SAM2 [16, 35] is a powerful vision foundation model, supports accurate video segmentation. To ensure consistency of 2D mask maps across views—i.e., that the same material region receives the same mask index from different views—we consider the multi-view image sequence as a video input and apply SAM2 for segmentation to generate the associated mask maps $\mathbf{M}=\{\mathbf{M}_{o}\}_{o=1}^{N}$.
@@
where $\mathbf{K}$ and $\mathbf{R}_{o}|\mathbf{T}_{o}]$ represent the camera’s intrinsic and extrinsic parameters, respectively. We use the 3DGS-estimated depth to check if the Gaussian kernel is visible in the segmentation map $\mathbf{M}_{o}$. If it is visible, we include the material group from this view in the voting process. After processing the segmentation maps for all views, we perform majority voting, assigning the material group $g_{k}$ that appears most frequently across all views to Gaussian kernel $\boldsymbol{\mathcal{G}_{k}}={(\mathbf{x}_{k},\mathbf{\Sigma}_{k},\alpha_{k},\mathbf{c}_{k},g_{k})}$. Repeating these steps allows us to determine the material groups for all Gaussian kernels $\mathcal{G}$.
@@ 3.3 Physical Internal Structure Discovery
Since the internal structure of an object can only be inferred from surface information available in the input image, we establish a surface-to-interior correspondence by assigning material groups from surface particles to internal ones. To ensure closure of the boundary set, we additionally designate the outermost layer of grid particles as surface particles. Each surface particle is then assigned to its nearest Gaussian kernel via a nearest-neighbor search and inherits its material group accordingly. For each internal particle, rays are cast along the six principal axes to collect material groups from the intersected surface particles. The most frequently observed material group $g_{i}$ is then assigned to the internal particle. Subsequently, we merge 3D Gaussians $\mathcal{G}$ with the filled internal particles $\mathcal{P}_{I}$ to obtain a unified continuum particles representation with material groups: $\mathbb{P}=\{\mathcal{G}|\mathcal{P}_{I}\}=\{\{(\mathbf{x}_{k},g_{k})\}|\{(\mathbf{x}_{i},g_{i})\}\}$. Based on this representation, we can easily assign material properties to different material regions of the 3D object.
@@ 3.4.1 Material Information Reasoning
In the real world, objects are typically composed of different materials. In 4D generation, users often lack the necessary physical knowledge to provide reasonable material properties for simulation. This greatly limits the physical realism of the simulation results. Recently, multimodal large language models have advanced rapidly, exhibiting knowledge far beyond that of humans, including rich physical prior knowledge. Inspired by this, we introduce GPT-4o [1], which reasons the material properties (e.g., Density $\rho$, Young’s modulus $E$, Poisson’s ratio $\nu$) of the internal and external parts of objects through vision.
Specifically, the user-input image is treated as the canonical view $\mathbf{I}_{c}$ for reasoning with the MLLMs. Sub-images representing different material components are extracted from this view based on the segmentation mask map. Following this, the canonical view and its segmented sub-images are fed into GPT-4o, prompting it to reason the internal and external material types and properties for the object described in each segmented sub-image. Detailed prompts are provided in Appendix A.2. The output format is as follows: {partial object 1: {surface: {material type, material property}, internal: {material type, material property}}, …}.
3.4.2 CLIP Fusion.
@@
Next, we calculate the similarity between the CLIP features of the canonical view and those of the rendered image. By selecting the matching pairs with the highest similarity, we establish a correspondence between the material groupings. Based on this, we can assign the surface and internal material information to the corresponding continuum particles. Finally, we construct a complete material continuum representation $\tilde{\mathbb{P}}={(\mathbf{x}_{p},g_{p},\theta_{p})}$, where $\theta_{p}$ denotes the material properties, for use in subsequent physical simulation to generate 4D dynamics.
@@
Phys4DGen can integrate any particle-based physical simulation algorithm. In this paper, we use the Material Point Method (MPM) [48] to simulate the dynamics of 4D content, which enables the modeling of motion and deformation behavior of continuum under external forces. For details on MPM and external forces, please refer to Appendix C. For physical simulation, we further assign temporal properties $t$ to the material continuum, along with other physical attributes involved in the simulation process, such as mass $m$, deformation gradient $\mathbf{F}$, and velocity $\mathbf{v}$. Then, we employ MPM to perform physical simulations on the material continuum $\tilde{\mathbb{P}}^{t}$. This allows us to track the position and local deformation of each particle at every time step:
@@
where $\mathbf{x}^{t+1}=\{\mathbf{x}_{p}^{t+1}\}_{p=1}^{P}$ denotes the positions of all particles at time step $t+1$. $\mathbf{F}^{t+1}=\{\mathbf{F}_{p}^{t+1}\}_{p=1}^{P}$ represents deformation gradients, which describe the local deformation of each particle at time step $t+1$. To reconstruct the 3D Gaussian Splatting (3DGS) representation at time step $t$ from the simulation results, we isolate the 3DGS-relevant components from the simulated material continuum. To incorporate the local deformation behavior of each GS kernel, we interpret the deformation gradient as a local affine transformation applied to the Gaussian kernel. Consequently, we can derive the covariance matrix of the Gaussian kernel $k$ in step $t+1$:
@@
At each step of the MPM simulation, we obtain the deformed 3DGS representation. The sequence of 3DGS representations across all time steps collectively forms the 4D content. This enables the generation of physically plausible 4D dynamics. Furthermore, since material grouping for 3D objects has already been established in Sec. 3.2, external forces can be precisely applied to any specific material region without manual selection, enabling user-friendly interaction and fine-grained control.